Nitrogen dioxide reacts with water to produce oxygen and ammonia. Using the ideal gas law equation, calculate the grams of NH3 that can be produced when 4.5p L of NO2 react at a temperature of 415 degrees Celsius and a pressure of 715mm Hg.
1
Expert's answer
2019-03-25T07:31:59-0400

Solution.

2NO2+3H2O=2NH3+2.5O2

pV=nRT

n(NO2)=pV/RT

n(NO2)=95325.5*4.5/8.31*688=75.03 mole

n(NO2)=n(NH3)

m(NH3)=n(NO2)*M(NH3)

m(NH3)=75.03*14=1275.51 g

Answer:

m(NH3)=75.03*14=1275.51 g